OpenStudy (anonymous):

A copy machine makes28 copies per minute. How long does it take to make 133copies? i need in minutes and seconds

OpenStudy (anonymous):

solve the proportion: \(\Huge \frac{28\text { copies}}{1\text { minute}}=\frac{133\text { copies}}{x\text { minutes}} \) can you find x first?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

why the unhappy face?

OpenStudy (anonymous):

i don't know the answer>>

OpenStudy (anonymous):

cross-multiply to get the equation: 28x = 133 can you solve this equation? :) or :( ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

28x = 133 x = 133/28 x represents the number of minutes it takes to make 133 copies.... use your calculator... x = ????

OpenStudy (anonymous):

4.75

OpenStudy (anonymous):

so 4.75 minutes.... how many seconds is .75 of a minute ???

OpenStudy (anonymous):

45